# Giant electrode effect on tunneling magnetoresistance and electroresistance in van der Waals intrinsic multiferroic tunnel junctions

### v1

@@ B. Significant TMR/TER and biaxial strain effects at equilibrium
Previous studies have suggested that applying in-plane biaxial strain is an effective measure to enhance the transport performance of vdW MFTJs [26, 31]. Considering the mismatch at the interfaces of our MFTJs, we systematically investigate the influence of in-plane biaxial strain in the range of -3% to 3% with a 1% interval on these MFTJs in the equilibrium state, and the results are presented in Fig. 6. It can be observed that biaxial strain significantly modulates TMR, TER, SIE, and RA. For Ag-VS-Au MFTJ, as shown in Fig. 6(a), regardless of P or P states, TMR increases with increasing compressive strain, and the TMR value changes from negative to positive. Conversely, tensile strain has a minor effect, maintaining negative values throughout. In contrast to TMR, overall TER decreases with increasing strain. From Fig. 6(b), it can be observed that in the PC state, the SIE of P/P remains around 80% unaffected by tensile strain, but decreases with increasing compressive strain. Figure 6(c) indicates that the RA of the MFTJ in the quadruple resistance state generally decreases during the transition from high tensile strain to compressive strain. For MoS-VS-PtTe MFTJ, as displayed in Fig. 6(d), only at a strain of -3%, the TMR becomes positive, while under other strains, it remains negative. Meanwhile, the TER reaches its maximum value at this compressive strain (-3%), approximately 200%. Furthermore, the SIE exhibits oscillatory behavior, peaking at 80% under a strain of -3% [see Fig. 6(e)]. Interestingly, as shown in Fig. 6(f), the RA for all four resistance states peaks at over 3 ·m under a strain of -1%. Above or below this strain value, the RA rapidly decreases to a minimum of approximately 0.1 ·m. Exciting stain effects are observed in the FeGaTe-VS-FeGeTe MFTJ. As displayed in Fig. 6(g), both TMR values for the two ferroelectric polarization states (P/P) increase with increasing compressive strain, while tensile stress remains unchanged. The strain has little effect on TER, with small numerical oscillations. Fig. 6(h) reflects that the SIE under the PC state remains constant at 100% within the studied strain range, indicating the robustness of the perfect spin filtering effect against strain. Like the Ag-VS-Au MFTJ, RA gradually decreases to 0.16 ·m during the transition from maximum tensile to compressive strain. Therefore, the aforementioned results indicate that strain serves as an effective approach for modulating the transport properties of MFTJs.
@@ C. Voltage-tunable transport properties in nonequilibrium state
Next, as shown in Fig. 7, we calculate the bias voltage-dependent (ranging from -0.5 V to 0.5 V) spin polarization current, spin injection efficiency (SIE), TMR ratio, TER ratio, and RA of these tunnel junctions in the PC(M)/APC(M) and P/P states. Note that the bias voltage, denoted as , is set by applying the chemical potential on the left (right) electrode as (). For the Ag-VS-Au MFTJ, as depicted in Fig. 7(a), it is evident that in the PC state, regardless of the ferroelectric polarization orientation, the current monotonically increases with increasing bias voltage, and the total current is mostly contributed by the spin-down channel [with the SIE greater than 0.8 in Fig. 7(c)], indicating a significant spin polarization rate. Under the APC state, the current exhibits a monotonic increase, but the rate of increase is slower and smaller compared to the PC state [see Fig. 7(b)]. As depicted in Fig. 7(d), one can observe that the TMR increases with bias voltage in both P and P states, and TMR becomes positive for bias voltages exceeding 0.2 V. The maximum TMR is less than 50%, and the bias voltage has a relatively minor impact on TER. Excitingly, as shown in Fig. 7(e), despite the increase in the resistance-area product (RA) with increasing bias voltage across the four resistance states, the maximum value remains below 0.06 ·m. Such a small value is beneficial for device performance.
Let us turn to the MoS-VS-PtTe MFTJ. As shown in Figs 7(f) and (g), the current-voltage (IV) characteristics of the PC and APC states resemble a diode-like behavior, where the diode conducts under positive bias voltage and exhibits reverse bias leakage current. At the equilibrium state, the SIE of this MFTJ is only around 0.5 (refer to Table 1). However, when the bias voltage is set to -0.2 V, it can be increased to above 0.8 as shown in Fig 7(h). It is evident that bias voltage has an adverse effect on both TMR and TER, with the maximum values occurring at zero bias voltage [see Fig 7(i)]. Interestingly, iFig 7(j), the RA across the four resistance states shows synchronized evolution under the different bias voltages. That is, it monotonically increases with positive bias voltage, reaching a minimum value of 0.11 ·m, while under negative bias voltage, it initially decreases and then increases, indicating the beneficial effect of positive bias voltage on enhancing the performance of this tunnel junction.
@@
The most significant bias voltage effect occurs in the FeGaTe-VS-FeGeTe MFTJ. In the PC state, whether it is P or P state, the current exhibits oscillatory behavior with increasing bias voltage, and the total current is contributed by the spin up current, implying negative differential resistance and perfect spin filtering effect [see Fig. 7(k)]. Compared to the PC state, as shown in Fig. 7(l), the current in the APC state remains nearly zero within the bias voltage range of -0.2 V to 0.2 V, indicating the presence of a giant magnetoresistance effect. As shown in Fig. 7(m), one can find that the SIE of the P/P under the PC state exhibits robustness against bias voltage within the range of -0.5 V to 0.3 V and maintains a value close to 1. This implies that the perfect spin filtering effect is not constrained by bias voltage, indicating strong controllability of spin current. More significantly, the bias voltage can markedly improve TMR and TER, as shown in Fig. 7(n). In both ferroelectric polarization P/P states, the TMR decreases with increasing negative bias voltage. However, it reaches its maximum (7600%/34000%) when the positive bias voltage increases to 0.1 V, after which it decreases rapidly. Under the PC state, the TER remains almost unaffected by the bias voltage, maintaining a constant value. However, under the APC state, the TER follows a similar trend to the evolution of TMR under both ferroelectric polarizations. The maximum TER occurs at 0.1 V and can reach up to 380%. In addition, the application of bias voltage in Fig. 7(o) is obviously detrimental to RA, but fortunately, it is less than 1 ·m in the range of -0.1 V to 0.1 V.
@@
Based on the above discussion, we can obtain that the FeGaTe-VS-FeGeTe MFTJ, under bias voltage, exhibits the optimal performance among tunnel junction devices due to its combination of maximum TMR/TER, perfect spin filtering, and negative differential resistance effects. Therefore, in comparison to the other two types of electrodes, we consider the van der Waals magnetic electrode to be the best choice.
D. Graphene intercalation effect in Ag-Gra-VS-Gra-Au MFTJ
@@
For the Ag-VS-Au MFTJ, in accordance with the aforementioned discussions, the strong hybridization between Ag/Au and adjacent VS weakens the charge redistribution ferroelectricity between the bilayer VS and alters the intrinsic electronic structure, ultimately resulting in its minimal TMR and TER. To preserve the intrinsic multiferroicity of bilayer VS, we introduce monolayer graphene (Gra) interlayers between Ag/Au and VS, referring to previous studies [31, 60]. The geometric structure of the new MFTJ device is denoted as Ag-Gra-VS-Gra-Au as shown in Figs. 8(d) and (e). In this new MFTJ, there are three different interfaces: Ag-Gra, Au-Gra, and VS-Gra. To determine the stacking configurations of these interfaces, we calculate the evolution of their total energy with stacking orders and interlayer distance, as shown in Figs. 8(a)-(c). From the inserted structure diagram in Fig. 8(a), one can observe that the VS-Gra heterostructure exhibits two sets of three equivalent stacking sequences: one comprising S-hollow/bond and V-C, and the other consisting of V-hollow/bond and S-C. The calculation results of the total energy indicate that the first combination (V-C) is the optimal stacking configuration for VS-Gra interface. For the Ag/Au-Gra interface, we calculate the evolution of the total energy with the interlayer distance for two stacking orders, as shown in Figs. 8(b) and (c). It can be observed that the optimal stacking order and interlayer distance are Ag/Au-C and 3.36/3.37 Å, respectively. The PDOS of the central scattering region in real space along the transport direction can clearly reflect the effect of graphene intercalation, as depicted in Figs. 8(f) and (g). Note that here we are only presenting the PDOS of the PC states under equilibrium conditions. From the spin up PDOS, it can be observed that a broad blue-black region emerges between the metal electrodes Ag/Au and VS, indicating a clear isolation of the wave function coupling between the two. Additionally, one can observe a blue-black region between the bilayer VS at the Fermi level. Contrasting with the mentioned above PDOS of Ag-VS-Au MFTJ, this reveals that graphene intercalation preserves the intrinsic properties of the bilayer VS.
@@
Table 2 summarizes the transport properties of the Ag-Gra-VS-Gra-Au MFTJ MFTJ at the equilibrium state. Clearly, under the influence of graphene intercalation, the values of all four non-volatile resistive states have been elevated by an order of magnitude. We also investigate the transport characteristics of this MFTJ under the influence of applied biaxial stress and bias voltage, as shown in Fig. 9. The TMR of both ferroelectric polarization states increases with increasing compressive stress (ignoring sign) and decreases with increasing tensile stress [see Fig. 9]. In general, apart from a 2% tensile stress, all other strain conditions have a weakening effect on TER. As depicted in Fig. 9(b), tensile strain can elevate the SIE close to 1 for the PC state, while it has almost no effect on the APC state. The smaller the RA of the MFTJ, the better its performance. It is evident that strain cannot significantly reduce RA compared to the case without strain [see Fig. 9(c)]. Figure 9(d) shows that under the PC state, the total current for both ferroelectric polarization arrangements P/P approaches linear increase with increasing bias voltage, with the majority contributed by the spin-down current. In the APC state, the current exhibits oscillatory behavior with bias voltage and surpasses the current under the PC state at certain biases, implying a negative TRM effect [refer to Fig. 9(e)]. As shown in Fig. 9(f), throughout the entire range of bias voltages, the SIE of the APC state is significantly smaller than that of the PC state. Specifically, the bias voltage can increase the maximum TMR to about 150%(-0.5 V), while the maximum TER occurs under non-biased conditions [see Fig. 9(g)]. For RA, as shown in Fig. 9(h), the RA of four resistive states increases under almost all bias voltages, indicating a detrimental effect on the performance of the MFTJ.
@@
Therefore, our analysis demonstrates that graphene intercalation is an effective strategy to enhance the transport performance of MFTJs and preserve the intrinsic multiferroic properties of bilayer VS. This finding could pave the way for novel experimental approaches.
@@ IV Summary
In summary, based on the first-principles density functional theory, we theoretically study the spin-dependent electronic transport properties of the bilayer VS-based vdW intrinsic MFTJs. Three different types of asymmetric electrode pairs are employed to investigate the influence on the transport properties of MFTJs. We demonstrate that the four giant non-volatile resistance states and resistance-area products (RA) in these MFTJs depend on electrode selection and the potential to manipulate these states by applying bias voltage and in-plane biaxial strain. At the equilibrium state, the maximum achievable TMR (TER) due to electrode effects is 10949% (153%), with the minimum RA being 0.026 ·m. Strain effects enhance TER to 208%, with RA decreasing to 0.016 ·m. Under non-equilibrium conditions, bias voltage further enlarges TMR (TER) to 34000% (380%). Comparing the three types of electrodes, we reveal that the electrode composition consisting of FeGaTe/FeGeTe offers the optimal choice for MFTJs due to the combination of giant TMR (TER), perfect spin filtering, and negative differential resistance effects. Furthermore, we also prove that graphene intercalation not only effectively eliminates the disruption of the hanging bonds in the Ag/Au electrodes for bilayer VS multiferroicity but also enhances the TMR (TER) by an order of magnitude. This work provides new design concepts for van der Waals intrinsic multiferroic tunnel junctions and paves the way for next-generation low-energy spintronics memory devices.
@@ V Acknowledgements
These authors contributed equally to this work.

### v2

@@ B. Significant TMR/TER and biaxial strain effects at equilibrium
Previous studies have suggested that applying in-plane biaxial strain is an effective measure to enhance the transport performance of vdW MFTJs [26, 31]. Considering the mismatch at the interfaces of our MFTJs, we systematically investigate the influence of in-plane biaxial strain in the range of -3% to 3% with a 1% interval on these MFTJs in the equilibrium state, and the results are presented in Fig. 6. It can be observed that biaxial strain significantly modulates TMR, TER, SIE, and RA. For Ag-VS2-Au MFTJ, as shown in Fig. 6(a), regardless of P or P states, TMR increases with increasing compressive strain, and the TMR value changes from negative to positive. Conversely, tensile strain has a minor effect, maintaining negative values throughout. In contrast to TMR, overall TER decreases with increasing strain. From Fig. 6(b), it can be observed that in the PC state, the SIE of P/P remains around 80% unaffected by tensile strain, but decreases with increasing compressive strain. Figure 6(c) indicates that the RA of the MFTJ in the quadruple resistance state generally decreases during the transition from high tensile strain to compressive strain. For MoS2-VS2-PtTe2 MFTJ, as displayed in Fig. 6(d), only at a strain of -3%, the TMR becomes positive, while under other strains, it remains negative. Meanwhile, the TER reaches its maximum value at this compressive strain (-3%), approximately 200%. Furthermore, the SIE exhibits oscillatory behavior, peaking at 80% under a strain of -3% [see Fig. 6(e)]. Interestingly, as shown in Fig. 6(f), the RA for all four resistance states peaks at over 3 ·m2 under a strain of -1%. Above or below this strain value, the RA rapidly decreases to a minimum of approximately 0.1 ·m2. Exciting stain effects are observed in the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ. As displayed in Fig. 6(g), both TMR values for the two ferroelectric polarization states (P/P) increase with increasing compressive strain, while tensile stress remains unchanged. The strain has little effect on TER, with small numerical oscillations. Fig. 6(h) reflects that the SIE under the PC state remains constant at 100% within the studied strain range, indicating the robustness of the perfect spin filtering effect against strain. Like the Ag-VS2-Au MFTJ, RA gradually decreases to 0.16 ·m2 during the transition from maximum tensile to compressive strain. Therefore, the aforementioned results indicate that strain serves as an effective approach for modulating the transport properties of MFTJs.
@@ C. Voltage-tunable transport properties in nonequilibrium state
Next, as shown in Fig. 7, we calculate the bias voltage-dependent (ranging from -0.5 V to 0.5 V) spin polarization current, spin injection efficiency (SIE), TMR ratio, TER ratio, and RA of these tunnel junctions in the PC(M)/APC(M) and P/P states. Note that the bias voltage, denoted as , is set by applying the chemical potential on the left (right) electrode as (). For the Ag-VS2-Au MFTJ, as depicted in Fig. 7(a), it is evident that in the PC state, regardless of the ferroelectric polarization orientation, the current monotonically increases with increasing bias voltage, and the total current is mostly contributed by the spin-down channel [with the SIE greater than 0.8 in Fig. 7(c)], indicating a significant spin polarization rate. Under the APC state, the current exhibits a monotonic increase, but the rate of increase is slower and smaller compared to the PC state [see Fig. 7(b)]. As depicted in Fig. 7(d), one can observe that the TMR increases with bias voltage in both P and P states, and TMR becomes positive for bias voltages exceeding 0.2 V. The maximum TMR is less than 50%, and the bias voltage has a relatively minor impact on TER. Excitingly, as shown in Fig. 7(e), despite the increase in the resistance-area product (RA) with increasing bias voltage across the four resistance states, the maximum value remains below 0.06 ·m2. Such a small value is beneficial for device performance.
Let us turn to the MoS2-VS2-PtTe2 MFTJ. As shown in Figs 7(f) and (g), the current-voltage (IV) characteristics of the PC and APC states resemble a diode-like behavior, where the diode conducts under positive bias voltage and exhibits reverse bias leakage current. At the equilibrium state, the SIE of this MFTJ is only around 0.5 (refer to Table 1). However, when the bias voltage is set to -0.2 V, it can be increased to above 0.8 as shown in Fig 7(h). It is evident that bias voltage has an adverse effect on both TMR and TER, with the maximum values occurring at zero bias voltage [see Fig 7(i)]. Interestingly, as shown in Fig 7(j), the RA across the four resistance states shows synchronized evolution under the different bias voltages. That is, it monotonically increases with positive bias voltage, reaching a minimum value of 0.11 ·m2, while under negative bias voltage, it initially decreases and then increases, indicating the beneficial effect of positive bias voltage on enhancing the performance of this tunnel junction.
@@
The most significant bias voltage effect occurs in the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ. In the PC state, whether it is P or P state, the current exhibits oscillatory behavior with increasing bias voltage, and the total current is contributed by the spin up current, implying negative differential resistance and perfect spin filtering effect [see Fig. 7(k)]. Compared to the PC state, as shown in Fig. 7(l), the current in the APC state remains nearly zero within the bias voltage range of -0.2 V to 0.2 V, indicating the presence of a giant magnetoresistance effect. As shown in Fig. 7(m), one can find that the SIE of the P/P under the PC state exhibits robustness against bias voltage within the range of -0.5 V to 0.3 V and maintains a value close to 1. This implies that the perfect spin filtering effect is not constrained by bias voltage, indicating strong controllability of spin current. More significantly, the bias voltage can markedly improve TMR and TER, as shown in Fig. 7(n). In both ferroelectric polarization P/P states, the TMR decreases with increasing negative bias voltage. However, it reaches its maximum (7600%/34000%) when the positive bias voltage increases to 0.1 V, after which it decreases rapidly. Under the PC state, the TER remains almost unaffected by the bias voltage, maintaining a constant value. However, under the APC state, the TER follows a similar trend to the evolution of TMR under both ferroelectric polarizations. The maximum TER occurs at 0.1 V and can reach up to 380%. In addition, the application of bias voltage in Fig. 7(o) is obviously detrimental to RA, but fortunately, it is less than 1 ·m2 in the range of -0.1 V to 0.1 V.
@@
Based on the above discussion, we can obtain that the Fe3GaTe2-VS2-Fe3GeTe2 MFTJ, under bias voltage, exhibits the optimal performance among tunnel junction devices due to its combination of maximum TMR/TER, perfect spin filtering, and negative differential resistance effects. Therefore, in comparison to the other two types of electrodes, we consider the van der Waals magnetic electrode to be the best choice. Note that considering lattice matching, we selected Fe3GaTe2/Fe3GeTe2 materials as magnetic electrodes, which inherently possess a high spin polarization. For MTJs/MFTJs, magnetic materials with high spin polarization will yield large TMR. Therefore, choosing magnetic materials with half-metal properties will result in even better transport properties.
D. Graphene intercalation effect in Ag-Gra-VS2-Gra-Au MFTJ
@@
For the Ag-VS2-Au MFTJ, in accordance with the aforementioned discussions, the strong hybridization between Ag/Au and adjacent VS2 weakens the charge redistribution ferroelectricity between the bilayer VS2 and alters the intrinsic electronic structure, ultimately resulting in its minimal TMR and TER. To preserve the intrinsic multiferroicity of bilayer VS2, we introduce monolayer graphene (Gra) interlayers between Ag/Au and VS2, referring to previous studies [31, 60]. The geometric structure of the new MFTJ device is denoted as Ag-Gra-VS2-Gra-Au as shown in Figs. 8(d) and (e). In this new MFTJ, there are three different interfaces: Ag-Gra, Au-Gra, and VS2-Gra. To determine the stacking configurations of these interfaces, we calculate the evolution of their total energy with stacking orders and interlayer distance, as shown in Figs. 8(a)-(c). From the inserted structure diagram in Fig. 8(a), one can observe that the VS2-Gra heterostructure exhibits two sets of three equivalent stacking sequences: one comprising S-hollow/bond and V-C, and the other consisting of V-hollow/bond and S-C. The calculation results of the total energy indicate that the first combination (V-C) is the optimal stacking configuration for VS2-Gra interface. For the Ag/Au-Gra interface, we calculate the evolution of the total energy with the interlayer distance for two stacking orders, as shown in Figs. 8(b) and (c). It can be observed that the optimal stacking order and interlayer distance are Ag/Au-C and 3.36/3.37 Å, respectively. The PDOS of the central scattering region in real space along the transport direction can clearly reflect the effect of graphene intercalation, as depicted in Figs. 8(f) and (g). Note that here we are only presenting the PDOS of the PC states under equilibrium conditions. From the spin up PDOS, it can be observed that a broad blue-black region emerges between the metal electrodes Ag/Au and VS2, indicating a clear isolation of the wave function coupling between the two. Additionally, one can observe a blue-black region between the bilayer VS2 at the Fermi level. Contrasting with the mentioned above PDOS of Ag-VS2-Au MFTJ, this reveals that graphene intercalation preserves the intrinsic properties of the bilayer VS2.
@@
Table 2 summarizes the transport properties of the Ag-Gra-VS2-Gra-Au MFTJ MFTJ at the equilibrium state. Clearly, under the influence of graphene intercalation, the values of all four non-volatile resistive states have been elevated by an order of magnitude. We also investigate the transport characteristics of this MFTJ under the influence of applied biaxial stress and bias voltage, as shown in Fig. 9. The TMR of both ferroelectric polarization states increases with increasing compressive stress (ignoring sign) and decreases with increasing tensile stress [see Fig. 9]. In general, apart from a 2% tensile stress, all other strain conditions have a weakening effect on TER. As depicted in Fig. 9(b), tensile strain can elevate the SIE close to 1 for the PC state, while it has almost no effect on the APC state. The smaller the RA of the MFTJ, the better its performance. It is evident that strain cannot significantly reduce RA compared to the case without strain [see Fig. 9(c)]. Figure 9(d) shows that under the PC state, the total current for both ferroelectric polarization arrangements P/P approaches linear increase with increasing bias voltage, with the majority contributed by the spin-down current. In the APC state, the current exhibits oscillatory behavior with bias voltage and surpasses the current under the PC state at certain biases, implying a negative TRM effect [refer to Fig. 9(e)]. As shown in Fig. 9(f), throughout the entire range of bias voltages, the SIE of the APC state is significantly smaller than that of the PC state. Specifically, the bias voltage can increase the maximum TMR to about 150%(-0.5 V), while the maximum TER occurs under non-biased conditions [see Fig. 9(g)]. For RA, as shown in Fig. 9(h), the RA of four resistive states increases under almost all bias voltages, indicating a detrimental effect on the performance of the MFTJ.
@@
Therefore, our analysis demonstrates that graphene intercalation is an effective strategy to enhance the transport performance of MFTJs and preserve the intrinsic multiferroic properties of bilayer VS2. This finding could pave the way for novel experimental approaches.
@@ IV Summary
In summary, based on the first-principles density functional theory, we theoretically study the spin-dependent electronic transport properties of the bilayer VS2-based vdW intrinsic MFTJs. Three different types of asymmetric electrode pairs are employed to investigate the influence on the transport properties of MFTJs. We demonstrate that the four giant non-volatile resistance states and low resistance-area products (RA) in these MFTJs depend on electrode selection and the potential to manipulate these states by applying bias voltage and in-plane biaxial strain. At the equilibrium state, the maximum achievable TMR (TER) due to electrode effects is 10949% (153%), with the minimum RA being 0.026 ·m2. Strain effects enhance TER to 208%, with RA decreasing to 0.016 ·m2. Under non-equilibrium conditions, bias voltage further enlarges TMR (TER) to 34000% (380%). Comparing the three types of electrodes, we reveal that the electrode composition consisting of Fe3GaTe2/Fe3GeTe2 offers the optimal choice for MFTJs due to the combination of giant TMR (TER), perfect spin filtering, and negative differential resistance effects. Furthermore, we also prove that graphene intercalation not only effectively eliminates the disruption of the hanging bonds in the Ag/Au electrodes for bilayer VS2 multiferroicity but also enhances the TMR (TER) by an order of magnitude. This work provides new design concepts for van der Waals intrinsic multiferroic tunnel junctions and paves the way for next-generation low-energy spintronics memory devices.